### v1

# A Greedy Quantum Route-Generation Algorithm

Abstract
Routing and scheduling problems with time windows have long been important optimization problems for logistics and planning. Many classical heuristics and exact methods exist for such problems. However, there are no satisfactory methods for generating routes using quantum computing (QC), for mainly two reasons: inequality constraints, and the trade-off of feasibility and solution quality. Inequality constraints are typically handled using slack variables; and feasible solutions are found by filtering samples. These challenges are amplified in the presence of noise inherent in QC. Here, we propose a greedy algorithm that generates routes by using information from all samples obtained from the quantum computer. By noticing the relationship between qubits in our formulation as a directed acyclic graph (DAG), we designed an algorithm that adaptively constructs a feasible solution. We prove its convergence to a feasible solution, and illustrate its efficacy by solving the Fleet Sizing Vehicle Routing Problem with Time Windows (FSVRPTW). Our computational results show that this method obtains a lower objective value than the current state-of-the-art annealing approaches, both classical and hybrid, for the same amount of time using D-Wave’s Hybrid Solvers. We also show its robustness to noise on D-Wave’s Advantage 4.1 through computational results as compared to the filtering approach on DWaveSampler, even when the filtering approach is given a longer annealing time, and a larger sample size.
Index Terms:
QUBO, quantum annealing, VRP, routing, greedy, D-Wave.
I Introduction
One of the most important problems in operations research is servicing customers within a certain time window. The simplest version of this is the travelling salesman problem with time windows, but more complex versions of this problem have also become relevant as transportation and logistics problems have become more complex. In this paper we study the Fleet Size Vehicle Routing Problem with Time Windows (FSVRPTW), which determines the minimum number of vehicles required to service all customers within their respective time windows.
The problem is practically-relevant on its own, but it is also a sub-problem in many solution methods for more complex routing problems such as column generation, bender’s decomposition, branch-and-price. However, in the aforementioned solution approaches, an exact solution method to the sub-problem is not always necessary - but a computationally-tractable one, is. Therefore, the community would benefit from faster solutions to this problem.
Quantum computing (QC) offers exciting possibilities to perform certain computational tasks better than classical solvers and may be able to speed up challenging optimization problems such as route planning. Certain algorithms such as Shor’s factorization or Grover’s search have proven advantage over their classical counterparts in fault-tolerant conditions. In the near future, it is not certain whether a quantum advantage can be observed due to the noise, and current size capabilities of quantum computers. As such, many hybrid approaches, involving classical and quantum computing, have been proposed. However, it still remains relevant to determine the best quantum approach to certain problems, so that when the QC capabilities become more advanced, the solution methods are readily available.
Routing problems on quantum computers have been solved since Feld et al [1]. However, no annealing-based methods have shown to be practical for generating routes themselves. Existing approaches that have shown to be capable of solving problems at scale generate routes classically, and then solve a set cover or exact cover problem. On the other hand, the existing annealing-based approaches do not scale well, due to slack variables, and other inefficient encoding schemes [2]. We review them briefly: [3] solves the traveling salesperson problem with time windows (TSPTW) using slack variables and one-hot encoding. This makes the number of qubits impractical. [4] uses position-based indexing. While this works under the single-agent setting in TSPTW, it loses much in optimality in a multi-agent setting because it assumes each agent makes a maximum number of stops. [5] creatively discretizes time using a state-based description, but the formulation of time-windows has a discrepancy with the capacity and state description. [6] presents several QUBO formulations but only solves small-sized problems, and, aside from the route-based formulation has difficulty obtaining feasible solutions. [7], [8] and [9] all present creative hybrid solution methods, but still generate routes classically.
The majority of existing QC solutions to Sherrington-Kirkpatrick (SK) Ising models involve sampling the QC a number of times, and eventually selecting the sample with the lowest energy. However, emerging research has shown the potential to use all of the samples from the QC, to take advantage of the correlation between samples [10]. We review them briefly here. [11] uses an iterative approach that fixes qubits based on their one-body expectation value, computed by using all of the samples returned from the annealer. Qubits with expectation values above a certain threshold are fixed to a particular value based on a majority vote, and a smaller SK Ising model is solved at the next iteration. The expectation value in some sense measures how certain we are of its value. [12], unlike [11], this paper has two separate procedures for selecting which qubits to fix, and the values they are fixed to. After selecting qubits, their values are determined by brute-force evaluation of the expectation value of the objective function, over all possible values of the selected variables. [13] uses two phases, instead of an iterative approach: 1) Run QAOA. 2) Compute the correlation matrix given by the output from sampling the QC circuit constructed from QAOA. The values of the principal eigenvector are then rounded to obtain a final solution. Other efforts to improve solution quality, e.g. unbalanced penalization [14], still cannot guarantee a feasible solution, and require heavy classical pre-processing for tuning penalty parameters. They also do not disclose computation time in their computational results.
However, these approaches do not directly apply to solve constrained optimization problems; the process of fixing variables may decrease the expectation value of the energy, but it might also violate constraints. The fixed values are maintained permanently throughout the remainder of the algorithm. For constrained problems, we require a delicate approach for ensuring a feasible solution can still be returned after the variables are fixed. We propose an algorithm which takes advantage of both the emerging research to utilize the correlation between samples, and also is guaranteed to return a feasible solution to the FSVRPTW. In this paper we present an annealing-based algorithm for generating routes under time window constraints that is suitable for the NISQ era [15] by its convergence to a feasible solution, and competitive computation time and solution quality. Our main contributions are as follows:
  • Present an annealing-based algorithm for generating routes, that represents the samples returned by the annealer as a DAG.
  • Prove that it converges to a feasible solution.
  • Show computational results on D-Wave, benchmarking it against classical, hybrid, and quantum annealing-based approaches based on computation time, and optimality, for a variety of practical-sized benchmark instances.
II Preliminaries
II-A Quadratic Unconstrained Binary Optimization
Many discrete optimization problems can be formulated as a Quadratic Unconstrained Binary Optimization (QUBO) problem. It is the common format used to input optimization problems into quantum annealers. Given square matrix $Q\in\mathbb{R}^{n\times n}$, the problem is to minimize $x^{T}Qx$ over binary variables $x_{i}\in\{0,1\}^{n}$:
$\min_{x\in\{0,1\}^{n}}x^{T}Qx$ (1)
Coefficients for linear terms in an optimization problem may be represented in the QUBO form by the diagonal elements of the matrix $Q$, since $x^{2}_{i}=x_{i}$. To represent this optimization problem on a quantum device, the QUBO is transformed into an Ising model with the Hamiltonian consisting of a weighted sum of tensor products of Pauli $Z$ operators. We use $Z_{i}$ to denote both the spin variables, and the Pauli $Z$ operators when clear from context. We may convert binary variables $x_{i}$ into spin variables $Z_{i}\in\{-1,1\}$ by $x_{i}=\frac{Z_{i}+1}{2}$ and represent them in an Ising model of the following Hamiltonian
$\mathcal{H}:=\sum_{i=1}v_{i}Z_{i}+\sum_{j<i}J_{ij}Z_{i}Z_{j}+C$ (2)
where $J_{ij},v_{i}$ encode the elements in the matrix $Q$. $C$ is a constant offset which is typically ignored since it doesn’t affect the optimal objective value. There is a one-to-one correspondence between the spin variables $Z_{i}$ and the binary variables $x_{i}$. Once a problem is cast as a QUBO, it may be solved on an annealer. A quantum annealer works by exploiting quantum mechanical effects to find low energy states of the Hamiltonian in equation (2) [16]. The output of an annealer is non-deterministic, so it is sampled many times to obtain a representative distribution. Given $M$ samples, the expected value of a variable $Z_{i}$ is $<Z_{i}>=\frac{\sum^{M}_{m}Z^{m}_{i}}{M}$, where $m$ is the index of the sample.
III Problem Formulation
We are given a timetable of customers, each with a time window $[e_{i},l_{i}]$, and a service time $q_{i}$; an example is shown in Table I. The start of service must being within the time window.
TABLE I: Timetable and distance matrix
Timetable
i $e_{i}$ $l_{i}$ $q_{i}$
0 0.0 $\infty$ 0.0
1 1.0 1.15 1.0
2 3.5 3.75 1.0
N 0.0 $\infty$ 0.0
Distance Matrix
i 0 1 2 N
0 0 1 3 0
1 1 0 2 1
2 3 2 0 3
N 0 1 3 0
We start by discretizing time such that for each time window, there exists a time discretization $t\in[e_{i},l_{i}]$. If an agent arrives at at customer before $e_{i}$, we allow for waiting. We create variables $x_{i,s,j,t}$ representing ”an agent leaves customer $i$ at time $s$, and subsequently leaves customer $j$ at time $t$”. We duplicate the depot to create original depot 00, and final depot $N$. The set $T$ is the set of time discretizations, the set $W$ is the set of customers, which excludes depots. For any variable $x_{i,s,j,t}$ we can compute $b_{ij}=\max\{e_{j},s+d_{ij}\}$ where $d_{ij}$ is the distance between customers $i$ and $j$. $b_{ij}$ is the earliest possible time we can start servicing customer $j$, having left customer $i$ at time $s$. We use an arc-based formulation from [6]: the constrained binary optimization (QUBO) is shown in equation (3).
The objective is in (3a). Constraint (3b) ensures that all customers are visited. Constraint (3c) enforces flow between consecutive customers. The remaining constraints are for pre-processing, and domain constraints.
$\min_{x_{i,s,j,t}}$ $\sum_{j\in W}\sum_{t}x_{0,0,j,t}$ (3a)
[…]
$\sum_{j\in\{0,W\},j\neq i,t}x_{j,t,i,s}=\sum_{j\in\{W,N\},j\neq i,t}x_{i,s,j,t},\,\forall i\in W,\forall s$ (3c)
$x_{i,s,0,t}=0,\,\forall s,t,\,\forall i\in\{0,W,N\}$ (3d)
$x_{N,s,j,t}=0,\,\forall s,t,\,\forall j\in\{0,W,N\}$ (3e)
$x_{0,s,j,t}=0,\,\forall s\neq 0,\forall t,\forall j\in\{0,W,N\}$ (3f)
$x_{i,s,j,t}=0,\,\forall i,s,t,\forall j\in\{0,W,N\}:b_{ij}>l_{j},$ (3g)
$x_{i,s,j,t}=0,\,\forall i,s,t,\forall j\in\{0,W,N\}:t<b_{ij}+q_{j}$ (3h)
$x_{i,s,i,t}=0,\,\forall s,t,\forall i\in\{0,W,N\}$ (3i)
$x_{i,0,j,t}=0,\,\forall i\neq 0,\forall t,\forall j\in\{0,W,N\}$ (3j)
$x_{i,s,j,0}=0,\,\forall i,s,\forall j\in\{0,W,N\}$ (3k)
$x_{0,s,N,t}=0,\,\forall s,t$ (3l)
$x_{i,s,j,t}\in\{0,1\}$ (3m)
III-A Pre-processing
Our formulation initially may appear to have a formidable number of variables. However, by making a few observations, we are able to prune many of the variables by fixing their values to 00, and removing them from the model. The variables we can prune correspond to constraints (3d)-(3l), and are described in Table II.
TABLE II: Constraint and Description for Pre-processing variables in problem (3)
Constraint Description
(3d) Cannot enter original depot
(3e) Cannot leave final depot
(3f) Can only leave original depot at time 0
(3g) Cannot be late to service customer $j$
(3h) Cannot leave customer $j$ before end of service time
(3i) No self-loops
(3j) Cannot leave customers at time 0
(3k) Cannot leave any nodes at time 0, after node $i$
(3l) Cannot travel immediately from original depot to final depot
IV Methods
IV-A Greedy Quantum Route-Finding Algorithm
Our algorithm for solving problem (3) is shown in Algorithm 1. We first observe that any subset of the variables in our formulation can be represented as a directed acyclic graph. By combining this with the samples obtained from the annealer, we can converge quickly to a feasible solution. The main idea, is the following: the annealer recommends a set of variables to be used in constructing a DAG, and a classical routine is used to find feasible sub-paths within the DAG. The variables corresponding to customers along each of the paths are pruned. Variables that haven’t been pruned are still ”active”. At each iteration the set of active variables $X$, gets smaller, while the cumulative set of feasible paths returned $S$, gets larger. The algorithm terminates when $X^{l}=\emptyset$, although in practice may be terminated when the number of active variables is small enough to solve the problem exactly.
We use $Q$ to denote the new set of edge-and-vertex-disjoint paths found in the DAG, and $P$ to denote a single path in $Q$. Paths consist of a sequence of tuples $\{(i^{P}_{k},s_{k})\}_{k=0}^{|P|}$. We use $i^{P}_{k}$ to denote the customer at index $k$ of path $P$, and no superscript is used when clear from context. Let $X^{l}$ denote the set of variables ”active” at iteration $l$, $S^{l}$ denote the cumulative set of paths found in iterations prior to $l$, and $Q^{l}$ the set of paths found at iteration $l$. We use (3b, $i$) to denote the coverage constraints for customer $i$, and (3c, $i,s$) to denote the flow constraints for customer $i$ at timestep $s$. DAG($X$) denotes the DAG formed by the variables in set $X$. QUBO($X$) denotes solving problem (3) using annealing, over the set of variables $X$. $\mathcal{X}$ denotes the samples returned by the annealer. $A,M$ are annealing time and number of samples, resp. They are implicitly inputs to QUBO($X$). When clear from context, $i$ is the index of a single variable in $X^{l}$. We use $R^{\frown}R^{\prime}$ to denote the concatenation of two paths $R$ and $R^{\prime}$. The final set of paths returned as a solution to the FSVRPTW (3) is $S$.
Algorithm 1 Greedy Quantum Route-Finding Algorithm
1:$X^{0}=\{x_{i,s,j,t}\}$ initially, all variables ”active”, $\theta$, $M$, number of annealing samples, $A$, annealing time.
2:$S,\text{paths found}$
3:$S^{0}\leftarrow\emptyset$
4:while $X^{l}\neq\emptyset$ do
5:     $\mathcal{X}^{l}\leftarrow$ QUBO($X^{l}$)
6:     for $i\in 1....|X^{l}|$ do
7:         Compute $|<Z_{i}>|$ over all $M$ samples
8:     end for
9:     $\bar{X}^{l},\leftarrow$ Select $\theta$ of $X^{l}$ with highest $|<Z_{i}>|$
10:     $Q^{l}\leftarrow$ Find paths in DAG($\bar{X}^{l}$)
11:     $S^{l+1},X^{l+1}\leftarrow$ Prune($S^{l},X^{l},Q^{l}$)
12:end while
Step 8 computes a statistic used in step 9 to determine which variables will participate in the DAG for finding paths. In our algorithm, we used one-body expectation values $|<Z_{i}>|$, although two-body expectation values may be used [12]. Parameter $\theta\in(0,1)$ is given by the user to determine the variables in $X^{l}$ to select at each iteration. $\theta$ may be given as a portion of the variables to select, or as a threshold for expectation values. In our running example and our analysis, we specify $\theta$ as a portion, although the results still hold for $\theta$ as a threshold. For step 10, any sub-routine may be used that guarantees that no 2 paths returned visit same customer. This ensures that the paths are edge-and-vertex-disjoint. Two examples are: returning the longest path, or returning multiple paths. Multiple paths may be returned by an iterative procedure: 1. Select the longest path in the DAG. 2. Remove all nodes (and their adjacent edges) in the DAG with the same customers $i_{k}$ (excluding depots $0,N$) along the currently-selected longest path, 3. Remove the nodes and arcs along the currently-selected longest path from the DAG. Repeat until the DAG is empty. In the simulation results, we returned multiple paths using this procedure. Heuristics such as returning a single longest path are more conservative, and may lead to longer computational time since fewer variables would be fixed at each iteration.
IV-B DAG Representation
We now describe the construction of the DAG. Any set of active variables $X^{l}$ may be represented as follows: A single node is created for each unique tuple $(i,s)$, where $i\in\{0,W,N\}$ and $s\in T$. A directed arc is present from $(i,s)$ to $(j,t)$ for each variable $x_{i,s,j,t}\in X^{l}$. By constraints (3g) and (3h), after the pre-processing described in section III-A variables only remain active if $s<t$. There cannot exist a cycle, otherwise it would imply that there exists a variable $x_{i,s,j,t}$ such that $s\geq t$. DAG($X^{0}$) for our running example is shown in Fig. 1.
IV-C Pruning
The subroutine Prune($S,X,Q$) is shown in Algorithm 2. Notice that the constraints (3b) and (3c) are separable by customer $j$, and tuples $(i,s)$ respectively. So we may prune variables and remove constraints independently for each tuple $(i,s)$, along a path in steps 6, 8, and 11 in Algorithm 2. If a path is found that shares an endpoint (either start or end) of a path in $S$, the paths are concatenated by Concat and the appropriate constraints are removed, as shown in Algorithm 3.
Algorithm 2 Prune
1:$S$, set of paths found previously, $Q$, new set of paths found.$X$, current set of active variables.
2:$S^{\prime}$, new set of paths found, $X^{\prime}$, new set of active variables, QUBO, new problem (3) formulation
3:for $P\in Q$ do
4:     QUBO $\leftarrow\textsc{Concat}(S,P,\text{QUBO})$
5:     for $(i_{k},s_{k})\in P$ do
6:         $X^{\prime}\leftarrow$ Apply pruning rules in section IV-C1.
7:         if $k\notin\{0,|P|\}$ then
8:              QUBO $\leftarrow$ remove constraint (3c,$i_{k},s_{k}$)
9:         end if
10:         if $k\neq 0$ then
11:              QUBO $\leftarrow$ remove constraint (3b,$i_{k}$)
12:         end if
13:     end for
14:end for
Algorithm 3 Concat
1:$S$, $P$, QUBO
2:QUBO, $S$
3:for $P^{\prime}\in S$ do
4:     if $i^{P}_{0}=i^{P^{\prime}}_{|P^{\prime}|}$ for $(i^{P}_{0},s)\in P$ then
5:         $P^{\prime}\leftarrow P^{\prime}\,{}^{\frown}P$
[…]
7:     end if
8:     if $i^{P}_{|P|}=i^{P^{\prime}}_{0}$ for $(i^{P}_{|P|},s)\in P$ then
9:         $P^{\prime}\leftarrow P^{\frown}P^{\prime}$
[…]
11:     end if
12:end for
Figure 1: Initial set of variables $X^{0}$ represented as a DAG (denoted DAG($X^{0}$)), continued from example in Tab. I, labeled with possible one-body expectation values from step 8 in Algorithm 1. Variables selected in step 9 in Algorithm 1, using $\theta=0.5$ are marked with solid arrows. Paths found in step 10 in Algorithm 1 are highlighted in blue.
Figure 2: Pruning example for $\theta=0.9$. Arcs in paths found in step 10 of Algorithm 1 are highlighted blue, and variables along the path are set to $1$ and pruned. Arcs highlighted red are set to 00 and pruned. Paths found are $P_{1}=\{(1,1),(N,2)\}$, $P_{2}=\{(0,0),(2,3),(N,3)\}$.
TABLE III: Pruning Rules Applied to Fig. 2
{tblr}
cell21 = r=6, cell22 = r=5, cell23 = r=5, cell81 = r=4, cell92 = r=2, cell93 = r=2, hline1,14 = -0.08em, hline2,8,12 = -, hline7,9,11 = 2-5, Path & $k$ $(i_{k},s_{k})$ Variables Pruned Pruning Rule
$P_{1}$ 00 $(1,1)$ $x_{1,1,N,2}=1$ Along path
$x_{0,0,1,2}=0$ exterior 1b
$x_{0,0,1,3}=0$ exterior 1b
$x_{1,3,N,3}=0$ exterior 1b
$x_{1,2,N,3}=0$ exterior 1b
$1$ $(N,2)$ None exterior 2a
$P_{2}$ 00 $(0,0)$ $x_{0,0,2,3}=1$ Along path
$1$ $(2,3)$ $x_{2,3,N,3}=1$ Along path
$x_{0,0,2,2}=0$ interior 2a
$2$ $(N,3)$ $x_{2,2,N,3}=0$ interior 1a
Figure 3: Subsequent iteration for pruning example with $\theta=0.9$. (Expectation values omitted to avoid clutter). After pruning in previous iteration, only variable $x_{0,0,1,1}$ remains active. Path $P_{3}=\{(0,0),(1,1)\}$ is found in step 10 of Algorithm 1, and joined with path $P_{1}$ in step 9 of Algorithm 3.
IV-C1 Pruning Rules
Throughout this section, the phrase ”prune” means the variable is set to the value specified and removed from the set of active variables $X^{l}$. For customer $i_{k}$ of tuple $(i_{k},s_{k})$, variables involving customer $i_{k}$ may be classified as:
  1. 1.
    $i=i_{k}$: ”outgoing” variables.
  2. 2.
    $j=i_{k}$: ”incoming” variables.
Index $k\in 0,...|P|$ along a path $P$ may be classified
  1. 1.
    $k\notin\{0,|P|\}$: ”interior” indices.
  2. 2.
    $k\in\{0,|P|\}$: ”exterior” indices.
Using this terminology we iterate over tuples $(i_{k},s_{k})$ and prune by the following rules:
Along path: If variable $x_{i_{k},s_{k},i_{k+1},s_{k+1}}$ is an arc along the path for two consecutive tuples $(i_{k},s_{k}),(i_{k+1},s_{k+1})$, it is set to $1$ and pruned.
If $k$ is ”interior”:
  1. 1.
    ”outgoing” variables are set to 00 and pruned if either:
    1. (a)
      $i=i_{k}$ and $s\neq s_{k}$ OR
    2. (b)
      $i=i_{k}$, and $s=s_{k}$, and either:
      1. i.
        $j\neq i_{k+1}$
      2. ii.
        $t\neq s_{k+1}$
  2. 2.
    ”incoming” variables are set to 00 and pruned if either:
    1. (a)
      $j=i_{k}$ and $t\neq s_{k}$ OR
    2. (b)
      $j=i_{k}$, and $t=s_{k}$, and either:
      1. i.
        $i\neq i_{k-1}$ OR
      2. ii.
        $t\neq s_{k-1}$
If $k$ is ”exterior”:
  1. 1.
    $k=0$:
    1. (a)
      $i_{k}=0$: Only prune and set $=1$ if $x_{i_{k},s_{k},i_{k+1},s_{k+1}}$ is Along path. (Variables incoming to depot were pruned by (3d)).
    2. (b)
      $i_{k}\neq 0$: Prune using rules ”interior” 1 and 2a.
  2. 2.
    $k=|P|$:
    1. (a)
      $i_{k}=N$: Prune using ”interior-incoming” 2a. (We do not prune according to rules 2b because incoming arcs to final depot $N$ at different times is allowed, as shown in Fig. 1).
    2. (b)
      $i_{k}\neq N$: Prune incoming variables under interior rule 2, and outgoing variables under interior rule 1a.
V Analysis
Lemma 1.
In any iteration $l$, if a customer $i^{P}_{k}$ is in the interior of path $P\in S^{l}$, all outgoing and incoming variables to that customer have been pruned.
Proof.
Assume without loss of generality that $P$ was formed by the union of paths $R$ and $R^{\prime}$, found at iterations $l^{R}$ and $l^{R^{\prime}}$, respectively. $P=R^{\frown}R^{\prime}$. For customer $i^{P}_{k}$ there are two cases:
  1. 1.
    $i^{P}_{k}$ was in the interior of either path $R$ or $R^{\prime}$. So it was either pruned to $1$ by the ”Along path” rules, or pruned to 00 by the ”interior” rules in section IV-C.
  2. 2.
    $i^{P}_{k}$ was in the exterior of of either path $R$ or $R^{\prime}$. The incoming variables to $i^{P}_{k}$ were pruned in iteration $l^{R}$ by ”exterior” rule 2b and the outgoing variables were pruned in iteration $l^{R^{\prime}}$ by ”exterior” rule 1b.
Lemma 2.
For every customer $i^{\prime}$, $\exists$ iteration $l^{\prime}$ in which the paths $Q^{l^{\prime}}$ returned in step 10 of Algorithm 1 will contain a path with a tuple $(i^{\prime},s)$ for some $s$.
Proof.
Every set of samples returned from the annealer $\mathcal{X}^{l}$ in iteration $l$ there are 2 cases:
  1. 1.
    Annealer returns 0 for all variables, for all samples. Then no variables are pruned, and the while loop resumes.
  2. 2.
    There exists a sample $\mathcal{X}^{l}_{m}\in\mathcal{X}^{l}$ with at least one variable $x_{i,s,j,t}=1$. Then at least one variable will be in $\bar{X}^{l}$ since $\theta\in(0,1)$. Then there are two cases:
    1. (a)
      $\bar{X}^{l}$ does not contain a variable $x_{i,s,j,t}$ with either $i^{\prime}=i$ or $i^{\prime}=j$. Variables containing customers along paths returned by step 10 will be pruned according to the pruning rules in section IV-C1, and the while loop resumes its next iteration with $X^{l+1}\subset X^{l}$.
    2. (b)
      $\bar{X}^{l}$ contains a variable $x_{i,s,j,t}$ with either $i^{\prime}=i$ or $i^{\prime}=j$. Then there are two cases:
      1. i.
        The paths returned by step 10 contain a tuple $(i^{\prime},s)$ for some $s$.
      2. ii.
        They do not.
In case 2(b)ii the variables containing customers along paths returned by step 10 will be pruned according to the pruning rules in section IV-C1, and the while loop resumes its next iteration with $X^{l+1}\subset X^{l}$. If in case 2(b)i. QED. If the annealing time $A$ is sufficiently large, [17] eventually we will exit case 1. If we continue to be in cases 2a or 2(b)ii above, then paths returned in step 10 in Algorithm 1 do not contain tuple $(i^{\prime},s)$, and in the next iteration $X^{l+1}\subset X^{l}$. By Lemma 1, all variables in the interior of previously-found paths $S$ will be pruned. At some iteration $l^{\prime}$ the only variables not pruned are those containing tuple $(i^{\prime},s)$, in which case they must appear in the paths returned in step 10. ∎
Lemma 3.
If customer $i^{R}_{k}$ was in the exterior of path $R\in Q^{l}$ in iteration $l$, then it will be in the interior of some path $P\in S^{l^{\prime}}$ for an iteration $l^{\prime}>l$.
Proof.
If $k=0$ then all outgoing variables and incoming variables to $i^{R}_{0}$ with $t\neq s^{R}_{0}$ have been pruned in ”exterior” rule 1b. So, the only active variables involving $i^{R}_{0}$ have $(j,t)=(i^{R}_{0},s^{R}_{0})$. By Lemma 2, in iteration $l^{\prime}$ path $R^{\prime}\in Q^{l^{\prime}}$ will contain $(i^{R}_{0},s^{R}_{0})$. This can only happen if $i^{R^{\prime}}_{|R^{\prime}|}=i^{R}_{0}$. In step 9 in Algorithm 3 these paths are unioned, and the resulting path $P=R^{{}^{\prime}\,{}^{\frown}}R$ will have $i^{R}_{0}$ in its interior.
The proof for $k=|R|$ is the similar, except $i^{R^{\prime}}_{0}=i^{R}_{|R|}$, and the resulting path is $P=R^{\frown}R^{\prime}$. ∎
Lemma 4.
Algorithm 1 converges to a feasible solution to Problem (3).
Proof.
Observe that all variables contain a customer by pre-processing constraint (3l). By Lemma 3 at some iteration, any customer $i$ will be in the interior of some path $P\in S$. By the pruning rules, when a customer is in the interior of a path $P\in S$ we have set exactly two variables involving customer $i$ equal to $1$: one incoming, and one outgoing. All other variables involving customer $i$ have been pruned and set to 00. Thus constraint (3b) is satisfied. By the same argument, if $(i,s)$ is along a path in $S$ returned by Algorithm 1 then exactly one variable on each side of constraint (3c,$i,s$) will be $1$ all others have been pruned by rules ”interior” 1 and 2. If $(i,s)$ is not along a path, then constraints (3c, $i,s$) will become $0=0$ and will be satisfied again by rules ”interior” 1 and 2. Since constraints (3b) and (3c) are satisfied the paths $S$ returned by Algorithm 1 are a feasible solution to Problem (3). ∎
VI Computational Results
VI-A Solvers
We benchmark with both classical and quantum annealing approaches that attempt to find a feasible solution to the FSVRPTW. In the absence of obvious heuristics for restoring feasibility, the standard approach is to run quantum annealer to obtain a distribution of samples, and select the feasible sample with the lowest energy. We give a brief description of each method used:Simulated annealing (SAS) is a classical metaheuristic algorithm that mimicks the physical process of annealing. Its algorithmic parameters include an annealing schedule. We use the SimulatedAnnealingSampler (SAS) provided by D-Wave, and we use the default settings of a geometric schedule [18]. LeapHybridCQMSampler (CQM) is a hybrid classical/quantum annealer that is specifically designed to solve problems with constraints. Its algorithmic parameters include num_reads, which is the number of samples returned by the annealer. The more samples, the greater chance of finding a solution with low energy. We used D-Wave’s CQM version 1p solver, and set num_reads=1000. In our greedy algorithm we used CQM for solving the QUBO($X$) in step 1, with the default parameters, and the default penalty values. To evaluate the performance of our proposed algorithm for practical-sized problems we used Solomon’s benchmark instances [19], briefly described as graphs with nodes that must be serviced with a time window, placed according to a random distribution. We used instance types $R101$/$R201$ which indicates narrow/wide time windows resp. We modify an instance to suit our problem as follows: for a given number of customers, randomly sample $N$ customers. For each value of $N$, we construct 10 random examples independently, and aggregate their statistics.
From Fig. 4(a), and 4(b) we can see that our greedy approach enjoys a small relative optimality gap as compared to SAS and CQM, in particular for large problem sizes $N>15$. This difference is even more pronounced in the R201 examples. For $N=25$, SAS returned feasible solutions for only 2 of the 10 random examples, and for $N=50$ it returned no feasible solutions. Statistics are aggregated only over feasible solutions. Fig. 5(a), and 5(b) show the relative time difference between the greedy algorithm, and CQM/SAS. Positive values indicate greedy algorithm is faster. For large problem instances greedy is always faster than SAS, although CQM is faster by a small margin for $N>15$.
(a) Solomon R101 Examples
(b) Solomon R201 Examples
Figure 4: Average relative optimality gap for SAS, CQM, and the proposed greedy algorithm, averaged over 10 examples with standard deviation. Relative optimal value is computed as $\frac{C_{A}-C_{opt}}{C_{opt}}$ where $C_{A}$ is the objective value found by the annealing method, and $C_{opt}$ is the optimal value found by Gurobi version 10.0.3. Statistics are computed only over feasible solutions. Data not shown for SAS, $N=50$, because no feasible solutions were found.
(a) Solomon R101 Examples
(b) Solomon R201 Examples
Figure 5: Average relative time difference of greedy as compared to SAS and CQM averaged over 10 examples. Relative time difference is computed as $\frac{t_{A}-t_{greedy}}{t_{A}}$ where $t_{A}$ is the time taken by the annealing method, and $t_{greedy}$ is the time taken by our proposed greedy algorithm. Statistics are computed only over feasible solutions. Data not shown for SAS, $N=50$, because no feasible solutions were found.
TABLE IV: QPU Results
$N$ DWaveSampler ($A=50\mu s$, $M=2000$) % Valid DWaveSampler ($A=50\mu s$, $M=2000$) with Pause % Valid $t_{\text{greedy}}$ $\alpha$ Avg. # logical Avg. # physical
5 20 10 0.08 10 44 415
6 0 0 0.074 17 84 601
7 0 0 0.08 33 109 890
8 0 0 0.09 0 121 955
9 0 0 0.09 0 127 1021
VII QPU Results
To evaluate its robustness to noise, our algorithm was also tested on DWave Advantage 4.1. For a fair benchmark, we also tried tuning the parameters of the DWaveSampler. An annealing pause was used by setting it to: [0.0, 0.0], [10.0, 0.4], [50.0, 0.4], [120, 1.0]. Annealing pauses are showng to perform well for scheduling problems [20]. Finally, to increase the chance of the DWaveSampler finding a feasible solution, both the annealing time, and the number of samples were adjusted to $50\mu s$ and $M=2000$, respectively. (This is the largest these values can be increased before the DWave’s QPU access time is exceeded [21]). In our greedy algorithm, we left the annealing parameters in DWaveSampler to their default values.
Our experiments revealed that even after enhancing the DWaveSampler with these features, it fails to find any feasible solutions in any of the 10 examples for problem sizes greater than N=5 customers. The greedy algorithm, when using the default DWaveSampler parameters, still maintains an average relative optimality gap below 33%. The total time of Algorithm 1, and the average relative optimality gap are labeled $t_{\text{greedy}}$, and $\alpha$, respectively, Table IV.
To show the scaling of this formulation on the binary variables, the average number of logical qubits and physical qubits are also reported. When the number of customers exceeds 7, the rate at which the number of logical qubits increases more slowly. This can be explained by the fact that only one discretization value within each time window is needed. If there is already one time discretization within a time window, then we avoid creating a new one. As a timetable includes more customers, more of the time windows overlap.
Conclusions and Future Work
This work presents the first effective algorithm for generating routes using an annealing-based approach. We demonstrated its efficiency by solving the FSVRPTW. We prove that it converges to a feasible solution, and we benchmark it against state-of-the-art classical and hybrid annealing-based approaches using D-Wave. We showed it enjoys a smaller optimality gap than other annealing-based approaches, at a competitive computation time, even for practical-sized problems of 50 customers. It also shows to be noise-robust when executed on a QPU, by taking advantage of the entire sampleset returned by the annealer. Natural directions for future work include determining the dependence of the proposed algorithm’s performance on the dataset, e.g. using different time discretizations and graphs generated from different distributions. We may also explore the effect of the control parameters, e.g. $\theta$. These future directions will provide insight into how this adaptive algorithm may be used to solve a more general class of problems outside of routing and scheduling.
VIII Acknowledgements
J.M. acknowledges Ketan Savla, and David E. Bernal Neira for helpful input and discussion during the development of these ideas.